OpenStudy (anonymous):

Write the ratio as a unit rate. 286 miles in 5 and a half hour.

OpenStudy (hannahc234):

286 : 5.5 = (286/5.5) / (5.5)/(5.5) = 52/1

OpenStudy (anonymous):

Hey owletgot

OpenStudy (anonymous):